Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP Sample Environment PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 71     Entry time: 31 Mar 2024, 23:20
Author: Adam Round 
Group: SPB 
Subject: Shift end 

X-ray delivery

  • 12.72 keV, 10Hz, ~1.4 mJ/pulse, some beam drift in the horizontal direction observed (probably mirror related)

Optical laser delivery

  • 487 nm ns laser for pump-probe
  • 800 nm PP laser for jet illumination, stable delivery

Achievements/Observations

  • Collected PP (pump-probe) dataset with 5 us time delay more than 12 hours
  • Jet and data collection was stable most of the time
ELOG V3.1.4-7c3fd00